OpenStudy (anonymous):

How do I reduce 5*sqrt(3)/10*sqrt(5)?

OpenStudy (anonymous):

sqrt(3)/2*sqrt(5)? or sqrt(15)/10

OpenStudy (lgbasallote):

\[\frac{5\sqrt 3}{10\sqrt 5} = \frac{\sqrt 3}{2\sqrt 5}\] now rationalize by multiplying numerator and denominator by sqrt 5 \[\frac{\sqrt 3 \times \sqrt 5}{2\sqrt 5 \times \sqrt 5}\] \[\frac{\sqrt{15}}{2(5)} = \frac{\sqrt{15}}{2\sqrt{10}}\] got it @Leaper

OpenStudy (anonymous):

\[5\sqrt{3}/10\sqrt{5} = \sqrt{3}/2\sqrt{5}\] Now you can rationalise it by multiplying and diving by Square root of 5.. You will get : \[(\sqrt{3} \times \sqrt{5}) / (2\sqrt{5} \times \sqrt{5}) = \sqrt{15}/10\] Answer is : \[\sqrt{15 }/ 10\]

OpenStudy (anonymous):

@lgbasallote, there should come 10 in the denominator in the answer and not 2root10... Am I right in saying that???